Ria is 25 years younger than her mother. In 10 years from now her mother will be twice Ria's age. How old is Ria now?
I have tried a few equations but none of them have worked out so I think it would be best if you just help me from scratch :)
Thank you

Answer by richwmiller(17219)

r=m-25
m+10=2*(r+10)